# MixedGaussianAvatar: Realistically and Geometrically Accurate Head Avatar via Mixed 2D-3D Gaussian Splatting

### v2

@@ Introduction
$\mu_{g}^{2D}=\lambda R\mu_{l}^{2D}+T+p^{2D}_{\theta}$ (11)
where $T$ denotes the average position of the three vertices of the corresponding triangle in the global space. We introduced a learnable perturbation term $p^{2D}_{\theta}$ to adaptively correct positional errors.
@@
$\mu_{g}^{3D}=\lambda R\mu_{l}^{3D}+\mu_{g}^{2D}+p^{3D}_{\theta}$ (12)
Figure 4: Qualitative comparison on mesh reconstruction.
Progressive Training Strategy
2D Gaussian Training
@@
$L_{\text{rgb}}=(1-\lambda)L_{1}+\lambda L_{\text{D-SSIM}}$ (13)
@@
$L_{\text{pos}}=\left\|\max\left(\mu_{l}^{2D},\epsilon_{\text{pos}}\right)\right\|_{2}$ (14)
@@
$L_{\text{sca}}=\left\|\max\left(s_{l}^{2D},\epsilon_{\text{sca}}\right)\right\|_{2}$ (15)
To improve the performance of 2DGS in geometric modeling, we utilize their depth distortion function $L_{\text{d}}$ and normal consistency function $L_{\text{n}}$ (details can be found in  (Huang et al. 2024a)). Therefore, the training loss in this stage is expressed as:
$L_{\text{rec}}=L_{\text{c}}+\lambda_{1}L_{\text{pos}}+\lambda_{2}L_{\text{sca}}+\lambda_{3}L_{\text{d}}+\lambda_{4}L_{\text{n}}$ (16)
Mixed 2D-3D Gaussian Training
After completing training for 2DGS in Sec. 2D Gaussian Training and mixed 2D-3D Gaussian modeling in Sec. Mixed 2D-3D Gaussian Modeling, the scene now includes 2D Gaussians distributed across the surface of the geometry and 3D Gaussians created for color compensation. During the mixed 2D-3D Gaussian training stage, we train all parameters of the 3D Gaussians as well as the color parameters (spherical harmonics) of the 2D Gaussians. However, mixed 2D-3D Gaussians require that each 3D Gaussian is positioned close to its corresponding parent 2D Gaussian in global space. This proximity ensures that when both are projected into image space, they can significantly contribute to the same tile’s color. We propose a 2D-3D distance regularization term $L_{dis}$ that limits the distance between 2D and 3D Gaussians to a specified threshold. $\epsilon_{dis}$ is 1.
$L_{\text{dis}}=\left\|\max\left(\mu_{l}^{3D}+p_{\theta}^{3D},\epsilon_{\text{dis}}\right)\right\|_{2}$ (17)
@@
Experiments
Figure 5: Qualitative comparison of image rendering.
Experimental Settings
Dataset and Baselines
We use two challenging datasets: NeRSemble and INSTA. NeRSemble dataset (Kirschstein et al. 2023) is used to conduct reconstruction and rendering experiments on 9 subjects, with each recording including 16 viewpoints. To ensure a fair comparison, we use the same dataset split as GaussianAvatars. INSTA dataset (Zielonka, Bolkart, and Thies 2023) is used for rendering experiments. It consists of monocular videos from ten subjects, with the final 350 frames of each sequence set aside for testing.
For the NerSemble dataset, we selected NeRF-based methods, including PointAvatar (Zheng et al. 2023), and INSTA (Zielonka, Bolkart, and Thies 2023); and 3DGS-based methods, including FlashAvatar (Xiang et al. 2024), Gaussian Head Avatar (Xu et al. 2024), and GaussianAvatars (Qian et al. 2024). For the INSTA dataset, we selected NeRF-based methods, including NHA, IMAvatar (Zheng et al. 2022), and INSTA; and 3DGS-based methods, including FlashAvatar, Gaussian Head Avatar, and GaussianAvatars. Furthermore, we selected an industrial 3D reconstruction software, MetaShape (Agisoft 2023) as a baseline.
Evaluation Metrics
@@
Mesh Extraction and Texture Generation
To extract meshes from the reconstructed mixed Gaussians, we generate depth maps of the training views using the median depth values of the 2D Gaussians projected onto the pixels. To achieve a high-quality texture for the 3D head mesh, we utilize the RGB color from the mixed 2D-3D Gaussians projected onto the pixels. Then, we use Open3D (Zhou, Park, and Koltun 2018) to apply truncated signed distance fusion (TSDF) to merge the depth maps for mesh extraction and use the RGB color to generate the texture. For a fair comparison, we apply the same mesh extraction and texture generation method to the 3DGS-based baselines.
Comparison Results
@@
Table 1: Results on the NeRSemble (Kirschstein et al. 2023) and INSTA (Zielonka, Bolkart, and Thies 2022) dataset. Green indicates the best and yellow indicates the second.
@@
PointAvatar (Zheng et al. 2023) NeRSemble 0.0020 25.9 0.887 0.096
INSTA (Zielonka, Bolkart, and Thies 2022) 0.0018 26.8 0.895 0.117
FlashAvatar (Xiang et al. 2024) 0.0016 29.3 0.932 0.075
Gaussian Head Avatar (Xu et al. 2024) 0.0015 30.8 0.945 0.078
GaussianAvatars (Qian et al. 2024) 0.0013 31.4 0.941 0.064
(Ours) 0.0011 31.8 0.953 0.067
NHA (Grassal et al. 2022) INSTA 0.0024 27.0 0.942 0.043
IMAvatar (Zheng et al. 2022) 0.0021 27.9 0.943 0.061
INSTA (Zielonka, Bolkart, and Thies 2022) 0.0017 28.6 0.944 0.047
FlashAvatar (Xiang et al. 2024) 0.0017 29.2 0.949 0.040
Gaussian Head Avatar (Xu et al. 2024) 0.0016 29.7 0.958 0.043
GaussianAvatars (Qian et al. 2024) 0.0014 29.1 0.953 0.046
(Ours) 0.0012 30.4 0.962 0.039
Quantitative Evaluation
In this experiment, NeRSemble dataset was used to assess novel-view synthesis, while INSTA was used for evaluating self-reenactment. As shown in Tab. 1, it is clear that NeRF-based methods, such as PointAvatar and INSTA, generally yield lower image quality. In contrast, methods based on 3DGS tend to demonstrate stronger rendering capabilities. This observation highlights the inherent advantage of 3D Gaussians in color representation. FlashAvatar binds 3D Gaussians to the UV map, and controlling the number of Gaussians can lead to insufficient detail representation, resulting in lower rendering quality compared to other 3DGS-based methods. Tab. 2 presents the rendering results of our method on the NeRSemble dataset utilizing only 2D Gaussians. It is evident that, in comparison to methods based on 3DGS, the multi-view consistency of 2DGS greatly diminishes image quality. However, our method employs 3D Gaussians to overcome the limitations of 2D Gaussians in color rendering, resulting in higher-quality rendering than pure 2DGS. Our approach is highly competitive with 3DGS methods, but it significantly excels in geometric surface accuracy, outpacing 3DGS-based methods.
Qualitative Evaluation
@@
Table 2: Quantitative ablation study on NeRSemble dataset.
@@
Figure 6: Qualitative ablation study on the NeRSemble dataset.
Ablation Study
@@
Conclusion
In this work, we introduce MixedGaussianAvatar, a new method that employs mixed 2D-3D Gaussian Splatting for creating head avatars. We use 2DGS to maintain the surface geometry and employ 3DGS for color correction in areas where the rendering quality of 2DGS is insufficient, reconstructing a realistically and geometrically accurate 3D head avatar. We employ a local-to-global transformation technique to reconstruct dynamic head avatars and adopt a progressive training strategy to train mixed 2D-3D Gaussians. Our method demonstrates exceptional performance in both mesh reconstruction and texture rendering, setting a new standard in the field.

### v3

@@ 3.3.3. Local-to-Global Transformation
(11) $\mu_{g}^{2D}=\lambda R\mu_{l}^{2D}+T+p^{2D}_{\theta}(t)$
where $T$ denotes the average position of the three vertices of the corresponding triangle in the global space. We introduced a learnable perturbation term $p^{2D}_{\theta}(t)$ to adaptively correct positional errors, where $t$ represents the timestep.
@@
(12) $\mu_{g}^{3D}=\lambda R\mu_{l}^{3D}+\mu_{g}^{2D}+p^{3D}_{\theta}(t)$
Figure 4. Qualitative comparison on mesh reconstruction.
3.4. Progressive Training Strategy
3.4.1. 2D Gaussian Training
@@ 3.4. Progressive Training Strategy
(13) $L_{\text{rgb}}=(1-\lambda)L_{1}+\lambda L_{\text{D-SSIM}}$
@@
(14) $L_{\text{pos}}=\left\|\max\left(\mu_{l}^{2D},\epsilon_{\text{pos}}\right)\right\|_{2}$
@@
(15) $L_{\text{sca}}=\left\|\max\left(s_{l}^{2D},\epsilon_{\text{sca}}\right)\right\|_{2}$
To improve the performance of 2DGS in geometric modeling, we utilize their depth distortion function $L_{\text{d}}$ and normal consistency function $L_{\text{n}}$ (details can be found in  (Huang et al., 2024b)). Therefore, the training loss in this stage is expressed as:
(16) $L_{\text{rec}}=L_{\text{c}}+\lambda_{1}L_{\text{pos}}+\lambda_{2}L_{\text{sca}}+\lambda_{3}L_{\text{d}}+\lambda_{4}L_{\text{n}}$
3.4.2. Mixed 2D-3D Gaussian Training
After completing training for 2DGS in Sec. 3.4.1 and mixed 2D-3D Gaussian modeling in Sec. 3.3.1, the scene now includes 2D Gaussians distributed across the surface of the geometry and 3D Gaussians created for color compensation. During the mixed 2D-3D Gaussian training stage, we train all parameters of the 3D Gaussians as well as the color parameters (spherical harmonics) of the 2D Gaussians. However, mixed 2D-3D Gaussians require that each 3D Gaussian is positioned close to its corresponding parent 2D Gaussian in global space. This proximity ensures that when both are projected into image space, they can significantly contribute to the same tile’s color. We propose a 2D-3D distance regularization term $L_{dis}$ that limits the distance between 2D and 3D Gaussians to a specified threshold. $\epsilon_{dis}$ is 1.
(17) $L_{\text{dis}}=\left\|\max\left(\mu_{l}^{3D}+p_{\theta}^{3D},\epsilon_{\text{dis}}\right)\right\|_{2}$
@@ 3.4.2. Mixed 2D-3D Gaussian Training
4. Experiments
Figure 5. Qualitative comparison of image rendering.
4.1. Experimental Settings
4.1.1. Dataset and Baselines
We use two challenging datasets: NeRSemble and INSTA. NeRSemble dataset (Kirschstein et al., 2023) is used to conduct reconstruction and rendering experiments on 9 subjects, with each recording including 16 viewpoints. To ensure a fair comparison, we use the same dataset split as GaussianAvatars. INSTA dataset (Zielonka et al., 2023b) is used for rendering experiments. It consists of monocular videos from ten subjects, with the final 350 frames of each sequence set aside for testing.
For the NerSemble dataset, we selected NeRF-based and its variant methods, including PointAvatar (Zheng et al., 2023) and INSTA (Zielonka et al., 2023b); and 3DGS-based methods, including Gaussian Head Avatar (Xu et al., 2024) and GaussianAvatars (Qian et al., 2024a). For the INSTA dataset, we selected NeRF-based methods, including NHA, IMAvatar (Zheng et al., 2022), and INSTA; and 3DGS-based methods, including FlashAvatar, Gaussian Head Avatar, and GaussianAvatars. Additionally, we selected a 3D reconstruction industrial software, MetaShape (Agisoft, 2023) as a baseline.
4.1.2. Evaluation Metrics
@@
4.1.3. Mesh Extraction and Texture Generation
To extract meshes from the reconstructed mixed Gaussians, we generate depth maps of the training views using the median depth values of the 2D Gaussians projected onto the pixels. To achieve a high-quality texture for the 3D head mesh, we utilize the RGB color from the mixed 2D-3D Gaussians projected onto the pixels. Then, we use Open3D (Zhou et al., 2018) to apply truncated signed distance fusion (TSDF) to merge the depth maps for mesh extraction and use the RGB color to generate the texture. For a fair comparison, we apply the same mesh extraction and texture generation method to the 3DGS-based baselines.
4.1.4. CUDA Implementation
Figure 6. The CUDA implementation of mixed 2D-3D Gaussian Splatting.
Figure 7. The CUDA implementation of error-based selection.
We have implemented mixed 2D-3D Gaussian Splatting with CUDA kernels based on the 2DGS (Huang et al., 2024b) and 3DGS (Kerbl et al., 2023) frameworks.
In the implementation of error-based selection in mixed 2D-3D Gaussian modeling, we extract the point list and point list keys from the 2DGS CUDA rasterizer. These two lists are one-to-one correspondences, with each element corresponding to the projection of a Gaussian in the image space, and they are ordered by depth. Each element in the point list keys is a 64-bit integer, where the first 32 bits store the tile’s ID, and the last 32 bits store the depth value of the current Gaussian projection. Each element in the point list is a 32-bit integer, which stores the index of the 2D Gaussian corresponding to the Gaussian projection. Therefore, using these two lists, we can establish the mapping between the tiles and the 2D Gaussians. By using the tile’s ID, we can identify the 2D Gaussians that contribute to it, as shown in line 8 of Alg. 1. For example, as shown in Fig. 7, after the 2D Gaussian training process, we calculate that tiles 1 and 2 have the largest MSE. Based on the point list keys and point list, we have identified the corresponding 2D Gaussian indices as 235, 234, and 127.
In the implementation of mixed 2D-3D Gaussian splatting, we pass the connection between 2D and 3D Gaussians to the CUDA rasterizer. Fig. 6 is a schematic diagram of this process. During the forward pass, for each pixel, we perform alpha blending of the projections of all 2D Gaussians in the current tile in depth order, from front to back. When a 2D Gaussian with a 3D Gaussian child node is encountered, the projection of the 3D Gaussian is inserted after the projection of the current 2D Gaussian, and alpha blending continues. In the backward pass, this order is reversed, meaning that alpha blending is performed from back to front. Therefore, the projection of the child 3D Gaussian will be inserted in front of its parent 2D Gaussian projection, and the alpha blending of the child 3D Gaussian will be computed first. During the backward pass, we update the gradients of all parameters for the 3D Gaussians, as well as the gradients for the color parameters of the 2D Gaussians.
4.2. Comparison Results
@@
Table 1. Results on the NeRSemble (Kirschstein et al., 2023) and INSTA (Zielonka et al., 2022) dataset. Green indicates the best and yellow indicates the second.
@@
PointAvatar (Zheng et al., 2023) NeRSemble 0.0020 25.9 0.887 0.096
INSTA (Zielonka et al., 2022) 0.0018 26.8 0.895 0.117
[…]
NHA (Grassal et al., 2022) INSTA 0.0024 27.0 0.942 0.043
IMAvatar (Zheng et al., 2022) 0.0021 27.9 0.943 0.061
INSTA (Zielonka et al., 2022) 0.0017 28.6 0.944 0.047
[…]
4.2.1. Quantitative Evaluation
In this experiment, NeRSemble was used to assess novel-view synthesis, while INSTA dataset was used for evaluating self-reenactment. As shown in Tab. 1, it is clear that NeRF-based and its variant methods, such as PointAvatar and INSTA, generally yield lower image quality. In contrast, methods based on 3DGS tend to demonstrate stronger rendering capabilities. This observation highlights the inherent advantage of 3D Gaussians in color representation. FlashAvatar binds 3D Gaussians to the UV map, and controlling the number of Gaussians can lead to insufficient detail representation, resulting in relatively lower rendering quality compared to other 3DGS-based methods. “Ours w/o 3DGS” in Tab. 2 presents the rendering results of our method on the NeRSemble dataset utilizing only 2D Gaussians. It is evident that, in comparison to methods based on 3DGS, the multi-view consistency of 2DGS greatly diminishes image quality. However, our method employs 3D Gaussians to overcome the limitations of 2D Gaussians in color rendering, resulting in higher-quality rendering than pure 2DGS. Our approach is highly competitive with 3DGS methods, but it significantly excels in geometric surface accuracy, outpacing 3DGS-based methods.
4.2.2. Qualitative Evaluation
@@
Table 2. Quantitative ablation study on NeRSemble dataset.
@@
Figure 8. Qualitative ablation study on the NeRSemble dataset.
4.3. Ablation Study
@@
5. Conclusion
In this work, we introduce MixedGaussianAvatar, a new method that employs mixed 2D-3D Gaussian Splatting for creating head avatars. We use 2DGS to maintain the surface geometry and employ 3DGS for color correction in areas where the rendering quality of 2DGS is insufficient, reconstructing a realistically and geometrically accurate 3D head avatar. We employ a local-to-global transformation technique to reconstruct dynamic head avatars and adopt a progressive training strategy to train mixed 2D-3D Gaussians. Our method unifies image rendering and mesh reconstruction, using a mixed Gaussian representation to integrate the two modalities of 2D images and 3D meshes. It bridges multimedia expressions and multimodal interactions, seamlessly unifying 2D visual fidelity with 3D geometric authenticity for holistic avatar representation.